OpenStudy (anonymous):

is y-4 and 4-y are equal?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

no. take an example : if y = 1, y - 4 = 1 - 4 = -3 4 - y = 4 - 1 = 3

OpenStudy (anonymous):

they're opposed

OpenStudy (anonymous):

WHAT DO YOU MEAN?

OpenStudy (anonymous):

herm... i can't really explain more... do you know numbers with a - before ?

OpenStudy (anonymous):

negative numbers ?

OpenStudy (anonymous):

YEAH

OpenStudy (anonymous):

so, you see that -3 and 3 are differents ?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

ok thank you